a 0.01M solution of an acid HX has a pH of 2.5. what is the % ionization of HX?

Answers

A) 0%

B) 0.32%

C) 32%

D) 40%

E) 100%
Answer #1

Al+-0 At eam HA HA C 0 coca ca [H+] = ca. â om pH = 2.5 - log[nt] = 2.5 th:) = 162.5 [+) - 3.16x10 x= 3.16 *10 degree of ion